OpenStudy (erniellyd28):

x+y=6 x=2y ( solve the system of equations by using the substitution method)

OpenStudy (erak):

Basically just put the second equation into the first. x + y = 6 2y + y = 6 3y = 6 y = 2 Then solve for x by using that y. x + y = 6 x + 2 = 6 x = 6 - 2 x = 4

OpenStudy (erniellyd28):

so the answer is x=4

OpenStudy (erniellyd28):

?

OpenStudy (erak):

Your answer should be in (x,y) form

OpenStudy (erniellyd28):

there im lost sorry lol

OpenStudy (erak):

so the solution of a system of equations is just a point, written as (x,y). I solved for both x and y up there, just put it in the right form

OpenStudy (erniellyd28):

so would it be x=6 and y=2?

OpenStudy (erak):

well it's just (4,2)